What is the slope and y-intercept of 30x+5y=20
Thanks

Answer by mananth(16949)

30x+5y =20
5y= 20-30x
/5
y=4-6x
y=-6x+4
compare with y = mx+b
m=-6 the slope
b= 4 = y intercept